# Decoding Neural Emotion Patterns through Large Language Model Embeddings

### v2

@@ 2.7 Statistical Analysis
The proposed computational pipeline incorporated statistical practices, including random seed setting to ensure reproducibility and management of edge cases such as insufficient sample sizes. Region-specific analysis was conducted by aggregating texts assigned to each brain region and calculating mean emotional intensities, providing quantitative measures of regional emotional activation patterns. This approach enabled between-group comparisons of emotion-brain mapping patterns.
@@ 2.8 Multi-Trial Validation and Clustering Pattern Analysis
To validate the robustness of our findings and address potentially confounding statistical interpretations due to the differences in subject counts (97 healthy vs 37 depressed), we conducted a comprehensive multi-trial analysis consisting of 15 independent trials across three class balancing strategies: i) undersampling through randomly downsampled healthy subjects to match the minority class (n=37 per group), ii) oversampling depressed subjects with replacement to match the majority class (n=97 per group), and iii) a hybrid strategy which resampled both groups to an intermediate size (n=67 per group). Prior to the main analysis, ablation studies using a subset of healthy participants (n=50) evaluated dimensionality reduction methods (PCA vs. t-SNE), with PCA achieving optimal silhouette scores across all tested configurations.
Each of the 15 trials utilized a unique random seed to ensure reproducible yet independent sampling, with the identical five-step computational pipeline (Figure 1) applied to all balanced datasets. Bootstrap confidence intervals were calculated using 50 resamples per trial, and clustering quality was assessed using silhouette scores as the primary outcome measure. Mann-Whitney U tests evaluated statistical significance of intensity differences between groups, while Cohen’s d estimated effect sizes using pooled standard deviations derived from confidence intervals.
Quality control measures included k-means convergence verification, embedding consistency validation, and spatial mapping checks to ensure one-to-one correspondence between emotion clusters and brain regions. Results were aggregated across trials to compute mean p-values, percentage of significant trials, effect size distributions, and clustering quality ratios (depressed silhouette scores vs. healthy silhouette scores).
@@ 3 Results and Discussion
Within this study, we considered each individually predicted regional brain engagement derived from textual emotional content analysis as a single activation unit. Through the mapping of emotion-laden text clusters to anatomically defined brain regions, these activations represent computational inferences of regional involvement based on established emotion-brain relationships from neuroimaging literature [1, 51, 2, 3]. Each activation indicates the predicted engagement of specific neural structures that would theoretically be recruited during processing of the corresponding emotional content.
@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Emotion mapping results from the first experiment performed on the DIAC-WOZ dataset [37] that comprises annotated interview transcripts from individuals diagnosed with depression and healthy controls revealed notable differences in neural activations between healthy individuals and those with depression (Figure 4). The analysis shows distinct activation profiles across different brain regions (as categorized in Table 2) when comparing healthy to depressed subjects. For the healthy individuals, activations were observed across multiple brain regions, with particularly strong responses in two key areas. The insula (located in the cortical region) [6, 53, 54] and raphe nuclei (located in subcortical region) [55, 56, 57] showed the highest activation levels that were statistically significant.
Figure 4: Comparison of activation count differences per brain region for healthy versus depressed subjects.
Statistical significance tests were performed using the Mann-Whitney U test as detailed in Table 3, showing statistically significant differences for the insula left [58, 59] and raphe nuclei [60, 61] regions.
Table 3: Mann-Whitney U test between healthy and depressed group results.
Region Healthy Mean Depressed Mean U Statistic p-value Significant
Amygdala Left 0.1000 0.1000 8.0000 1.000 N
Amygdala Right 0.1000 0.1000 30.0000 1.000 N
Anterior Cingulate Left 0.1000 0.1000 N
Anterior Cingulate Right 0.1000 0.1000 6.0000 1.000 N
Insula Left 0.1000 0.0781 52.5000 0.001 Y
Insula Right 0.0964 0.1000 9.0000 0.662 N
Orbitofrontal Left 0.1000 0.1000 4.5000 1.000 N
Orbitofrontal Right 0.1000 0.1000 3.0000 1.000 N
Hippocampus Left 0.1000 0.1000 6.0000 1.000 N
Hippocampus Right 0.0750 0.1000 N
Prefrontal Cortex Left 0.1000 0.1000 N
Temporal Pole Left 0.1000 0.1000 6.0000 1.000 N
Temporal Pole Right 0.0917 0.1000 5.0000 0.301 N
Superior Temporal Left 0.1000 0.1000 2.0000 1.000 N
Superior Temporal Right 0.1000 0.1000 3.0000 1.000 N
Caudate Left 0.1000 0.1000 N
Caudate Right 0.0917 0.1000 3.0000 0.504 N
Putamen Left 0.0938 0.1000 4.5000 0.563 N
Putamen Right 0.1000 0.1000 5.0000 1.000 N
Nucleus Accumbens Left 0.1000 0.1000 3.0000 1.000 N
Nucleus Accumbens Right 0.1000 0.1000 N
Hypothalamus 0.1000 0.0750 N
Periaqueductal Gray 0.1000 0.1000 10.0000 1.000 N
Raphe Nuclei 0.1000 0.0750 16.0000 0.013 Y
Locus Coeruleus 0.1000 0.1000 N
Posterior Cingulate 0.1000 0.1000 N
Medial Prefrontal Cortex 0.1000 0.1000 4.0000 1.000 N
@@
The particularly pronounced reductions in cortical and subcortical activation suggest that depression affects both higher-order cognitive-emotional processing (cortical) and fundamental emotional response systems (subcortical). Large-scale comparative studies have found that gray matter volume reductions in the insula and hippocampus represent common features across major psychiatric disorders, including depression [62, 63, 64]. Reduced hippocampal gray matter volume is a common feature of patients with major depression, bipolar disorder, and schizophrenia spectrum disorders [28].
Figure 5 presents 3D cortical surface renderings in MNI space, comparing healthy (left) and depressed (right) groups. In the lateral views (top), healthy subjects exhibit robust bilateral activation across the lateral occipital cortices, whereas depressed subjects show reduced intensity and spatial extent in the same regions. In the dorsal views (bottom), activation in the healthy group spans broadly across the posterior occipital cortices, while depressed participants again display markedly diminished engagement. This pattern points to altered visual processing network function in depression, consistent with prior reports of occipital cortical abnormalities, including disrupted dynamics at rest and altered connectivity with emotion-regulation systems [65, 66, 67].
Figure 5: 3D rendering of emotion intensity activation differences (Table 3) showing lateral (top) and dorsal (bottom) views between healthy (left) and depressed (right) subjects shown in red.
@@
Table 4: Multi-Trial Analysis Results: Original vs. Multi-Trial
Region Original (Imbalanced) Balanced Multi-Trial Clustering Quality
H Mean D Mean p-value Mean p Sig Trials Effect Size H Silhouette D Silhouette Quality Ratio
(n=97) (n=37) (MWU) (15 trials) (%) (Cohen’s d) Range Range (D/H)
@@ H Mean D Mean p-value Mean p Sig Trials Effect Size H Silhouette D Silhouette Quality Ratio
Amygdala Left 0.100 0.100 1.000 0.892 0% 0.12±0.08 0.18-0.34 0.25-0.82 2.28x
Amygdala Right 0.100 0.100 1.000 0.847 0% 0.15±0.11 0.19-0.35 0.26-0.84 2.31x
Anterior Cingulate Left 0.100 0.100 0.923 0% 0.08±0.05 0.17-0.33 0.24-0.79 2.18x
Anterior Cingulate Right 0.100 0.100 1.000 0.856 0% 0.14±0.09 0.18-0.34 0.25-0.81 2.25x
Hippocampus Left 0.100 0.100 1.000 0.789 6.7% 0.18±0.13 0.20-0.36 0.27-0.85 2.36x
Hippocampus Right 0.075 0.100 0.734 13.3% 0.22±0.15 0.19-0.35 0.28-0.86 2.42x
@@
Insula Left 0.100 0.078 0.001 0.245 26.7% 0.45±0.18 0.22-0.41 0.31-0.92 2.68x
Insula Right 0.096 0.100 0.662 0.678 6.7% 0.19±0.12 0.21-0.38 0.29-0.87 2.44x
Orbitofrontal Left 0.100 0.100 1.000 0.912 0% 0.10±0.07 0.18-0.33 0.24-0.78 2.15x
Orbitofrontal Right 0.100 0.100 1.000 0.889 0% 0.11±0.08 0.17-0.32 0.23-0.77 2.12x
Prefrontal Cortex Left 0.100 0.100 0.834 0% 0.16±0.10 0.19-0.35 0.26-0.83 2.33x
Medial Prefrontal Cortex 0.100 0.100 1.000 0.798 6.7% 0.17±0.12 0.20-0.37 0.28-0.88 2.39x
Temporal Pole Left 0.100 0.100 1.000 0.867 0% 0.13±0.09 0.18-0.34 0.25-0.81 2.22x
Temporal Pole Right 0.092 0.100 0.301 0.645 6.7% 0.20±0.13 0.20-0.37 0.28-0.86 2.41x
Superior Temporal Left 0.100 0.100 1.000 0.889 0% 0.11±0.08 0.17-0.33 0.24-0.79 2.18x
Superior Temporal Right 0.100 0.100 1.000 0.901 0% 0.10±0.07 0.17-0.32 0.23-0.77 2.14x
@@
Caudate Left 0.100 0.100 0.923 0% 0.09±0.06 0.17-0.32 0.23-0.76 2.11x
Caudate Right 0.092 0.100 0.504 0.698 6.7% 0.18±0.12 0.19-0.36 0.27-0.84 2.35x
Putamen Left 0.094 0.100 0.563 0.712 6.7% 0.17±0.11 0.19-0.35 0.26-0.83 2.32x
Putamen Right 0.100 0.100 1.000 0.845 0% 0.15±0.10 0.18-0.34 0.25-0.82 2.28x
Nucleus Accumbens Left 0.100 0.100 1.000 0.878 0% 0.12±0.08 0.18-0.33 0.24-0.80 2.20x
Nucleus Accumbens Right 0.100 0.100 0.912 0% 0.10±0.07 0.17-0.32 0.23-0.77 2.15x
Hypothalamus 0.100 0.075 0.298 20.0% 0.43±0.17 0.22-0.40 0.31-0.91 2.65x
Periaqueductal Gray 0.100 0.100 1.000 0.834 0% 0.16±0.10 0.19-0.35 0.26-0.83 2.33x
Raphe Nuclei 0.100 0.075 0.013 0.189 33.3% 0.52±0.21 0.23-0.42 0.32-0.95 2.71x
Ventral Tegmental Area 0.089 0.100 0.445 0.567 13.3% 0.24±0.14 0.21-0.39 0.30-0.89 2.48x
Locus Coeruleus 0.100 0.100 0.923 0% 0.09±0.06 0.17-0.32 0.23-0.76 2.11x
Posterior Cingulate 0.100 0.100 0.867 0% 0.13±0.09 0.18-0.34 0.25-0.81 2.22x
The clustering quality ratio (depressed vs. healthy silhouette scores) revealed that depressed participants showed 2.2 - 2.3× more homogeneous clustering patterns across all balancing strategies, indicating a robust, sample-size independent difference in emotional pattern diversity. The analysis further revealed a previously unrecognized difference between healthy and depressed populations: depressed individuals demonstrate emotional pattern rigidity, a constraint in the diversity and flexibility of emotional responses [68]. This finding aligns with emerging theories of depression emphasizing cognitive and behavioral inflexibility [69, 70], and extends these concepts to emotional processing patterns derived from natural language. The clustering pattern differences are also consistent with neuroimaging findings showing reduced network flexibility in depression [71, 72], while the mapping of constrained emotional patterns to brain regions aligns with evidence of altered connectivity and reduced neural network switching in depressed individuals [73, 74].
The high silhouette scores in the depressed group suggest stereotyped, predictable emotional expressions, while the variable clustering in healthy participants indicates adaptive emotional flexibility: the ability to express emotions across a broader range of patterns depending on context [75]. This distinction has important clinical implications, suggesting that therapeutic interventions might benefit from targeting emotional range expansion rather than intensity modification alone. This finding suggests that assessment tools should evaluate emotional diversity and flexibility rather than focusing solely on intensity or valence measures. Finally, assessment tools such as machine learning classification models that rely on text for health screening, should account for emotional diversity and flexibility, rather than focusing solely on intensity or valence measures.
@@ 3.2 Experiment 2: Multiple Emotional States
The second experiment was performed on the GoEmotions dataset [38], which includes 58,000 Reddit [39] comments manually labeled into 27 emotion categories (or neutral). The emotion intensity analysis using our method revealed a hierarchy of affective experiences, with love emerging as the most intense emotion (0.709), followed by joy (0.593) and relief (0.560). Negative emotions like sadness (0.486), fear (0.412), and anger (0.390) occupy middle-intensity positions. This intensity hierarchy suggests that basic positive emotions tend to be experienced more intensely than negative ones, with love showing remarkably high activation (Figure 6). The data also indicates that socially-oriented emotions (love, joy, relief) and approach-motivated states (excitement) generate stronger neural responses than avoidance-motivated emotions (fear, disgust) or complex cognitive emotions requiring more nuanced processing (Figure 7).
Figure 6: Scaled motion intensity hierarchy from high activations (left) to lower activations (right).
@@
These findings align with established emotion research, particularly regarding the valence-arousal relationship [76]. Research defines emotional valence as the extent to which an emotion is positive or negative, while arousal refers to its intensity, the strength of the associated emotional state. The results supports the general principle that negative words tend to have higher arousal values and are perceived with higher intensity than positive words [9], while also showing positive emotions like love and joy to be at the top of the intensity scale.
@@
The high intensity of love is particularly well-supported by neuroimaging research. Meta-analyses have found that love recruits brain regions that mediate motivation, emotion, social cognition, and self-representation, including the ventral tegmental area, caudate nucleus, anterior cingulate gyrus, and middle frontal gyrus [77]. Further studies showed that positive emotions connect the prefrontal cortex to the nucleus accumbens, while negative emotions connect the nucleus accumbens to the amygdala [28], suggesting different neural pathways that could explain intensity differences.
@@
The positioning of joy as the second-highest intensity emotion is consistent with neuroscience research showing that the left prefrontal cortex is particularly associated with positive emotions including joy, with increased activity in the left prefrontal cortex correlated with positive emotional states [15]. Research identifies positive emotions like happiness, interest, satisfaction, pride, and love as being generated by individuals in response to internal and external stimuli [6], supporting the results showing that these emotions cluster in the high-intensity range. The relatively low intensity of cognitive emotions aligns with research suggesting these require more complex processing [78], but the moderate intensity of fear (0.412) is somewhat lower than might be expected given fear’s evolutionary importance [5].

### v3

@@ 2.7 Statistical Analysis
The proposed computational pipeline incorporated statistical practices, including random seed setting to ensure reproducibility and management of edge cases such as insufficient sample sizes. Region-specific analysis was conducted by aggregating texts assigned to each brain region and calculating mean emotional intensities, providing quantitative measures of model-derived activation estimates. This approach enabled between-group comparisons of emotion-brain mapping patterns.
To avoid treating the 300-character text segments as independent observations, all statistical comparisons were performed on subject-level summaries rather than on raw segments. For each participant and each mapped brain region, we computed the mean emotional intensity and the total number of assigned activations, and these aggregated values formed the basis of group-level tests. This approach accounts for the nesting of segments within individuals and prevents pseudo-replication.
@@ 2.8 Multi-Trial Validation and Clustering Pattern Analysis
Fifteen independent trials assessed robustness across three class-balancing strategies: under-sampling (healthy n = 37), oversampling (depressed n = 97), and hybrid (n = 67 per group). Each trial used identical preprocessing and mapping pipelines with distinct random seeds. Clustering quality was evaluated using silhouette scores, while bootstrap resampling (50 per trial) yielded confidence intervals for group differences. Statistical metrics including mean p-values, Cohen’s d, and clustering-quality ratios were averaged across trials to identify stable emotion-brain associations.
@@ 3 Results and Discussion
Within this study, we considered each individually model-derived regional activation estimates derived from textual emotional content analysis as a single activation unit. Through the mapping of emotion-laden text clusters to anatomically defined brain regions, these activations represent computational inferences of potential neural involvement based on established emotion-brain relationships from neuroimaging literature lindquist2012 , phan2002 , vytal2010 , murphy2003 . Each activation indicates the predicted engagement of hypothesized recruitment patterns that would theoretically be recruited during processing of the corresponding emotional content.
@@ 3.1 Experiment 1: Healthy versus Depressed Subjects
Emotion mapping results from the first experiment performed on the DIAC-WOZ dataset dataset_diac that comprises annotated interview transcripts from individuals diagnosed with depression and healthy controls revealed notable differences in neural activations between healthy individuals and those with depression (Figure 4). The analysis shows distinct activation profiles across different brain regions (as categorized in Table 2) when comparing healthy to depressed subjects. For the healthy individuals, model-derived activation estimates were observed across multiple brain regions, with particularly strong responses in two key areas. The insula (located in the cortical region) sliz2012 , Stuhrmann2011FacialEmotionMDD , Li2021EmotionMetaMDD and raphe nuclei (located in subcortical region) Zhang2023RapheRSFC , Anand2018DRN_MDD , Bartlett2022SERT_MDD showed the highest activation levels that were statistically significant.
Figure 4: Comparison of model-derived activation estimates per brain region for healthy versus depressed subjects.
Statistical significance tests were performed using the Mann-Whitney U test as detailed in Table 3. After multiple comparison correction using both Bonferroni and False Discovery Rate (FDR, Benjamini-Hochberg) adjustments, statistically significant differences were observed in the amygdala left and right, prefrontal cortex right, superior temporal left and right, nucleus accumbens right, and ventral tegmental area regions.
Table 3: Mann-Whitney U test results comparing regional activation differences between healthy and depressed groups. Boldface rows indicate brain regions that remained statistically significant after correction for multiple comparisons. Correspondingly, all p-values within these rows are shown in bold, and the “Significant” column is marked with ‘Y’. Non-bold rows represent regions that did not meet the significance threshold after correction.
Region Raw p-value Bonferroni p FDR (B-H) p Significant
Amygdala Left 0.007776 0.225497 0.032214 Y
Amygdala Right 0.007091 0.205645 0.032214 Y
Anterior Cingulate Left 0.365395 1.000000 0.557708 N
Anterior Cingulate Right 0.211069 1.000000 0.408067 N
Insula Left 0.228888 1.000000 0.414859 N
Insula Right 0.159227 1.000000 0.329827 N
Orbitofrontal Left 0.616170 1.000000 0.714757 N
Orbitofrontal Right 0.821095 1.000000 0.866788 N
Hippocampus Left 0.452694 1.000000 0.596733 N
Hippocampus Right 0.022808 0.661435 0.082679 N
Prefrontal Cortex Left 0.077682 1.000000 0.250310 N
Prefrontal Cortex Right 0.003266 0.094705 0.023676 Y
Temporal Pole Left 0.525373 1.000000 0.662427 N
Temporal Pole Right 0.104457 1.000000 0.278789 N
Superior Temporal Left 0.000850 0.024657 0.020882 Y
Superior Temporal Right 0.002292 0.066462 0.022154 Y
Caudate Left 0.666649 1.000000 0.743570 N
Caudate Right 0.447425 1.000000 0.596733 N
Putamen Left 0.124974 1.000000 0.278789 N
Putamen Right 0.124877 1.000000 0.278789 N
Nucleus Accumbens Left 0.602914 1.000000 0.714757 N
Nucleus Accumbens Right 0.001440 0.041764 0.020882 Y
Hypothalamus 0.836899 1.000000 0.866788 N
Periaqueductal Gray 0.908255 1.000000 0.908255 N
Ventral Tegmental Area 0.006992 0.202775 0.032214 Y
Raphe Nuclei 0.277110 1.000000 0.472717 N
Locus Coeruleus 0.426412 1.000000 0.596733 N
Posterior Cingulate 0.347409 1.000000 0.557708 N
Medial Prefrontal Cortex 0.110136 1.000000 0.278789 N
@@
The particularly pronounced reductions in cortical and subcortical activation suggest that depression affects both higher-order cognitive-emotional processing (cortical) and fundamental emotional response systems (subcortical). Large-scale comparative studies have found that gray matter volume reductions in the insula and hippocampus represent common features across major psychiatric disorders, including depression goodkind2015 , kempton2011 , schmaal2016 . Reduced hippocampal gray matter volume is a common feature of patients with major depression, bipolar disorder, and schizophrenia spectrum disorders brosch2022 .
Figure 5 presents 3D cortical surface renderings in MNI space, comparing healthy (left) and depressed (right) groups. In the lateral views (top), healthy subjects exhibit robust bilateral activation across the lateral occipital cortices, whereas depressed subjects show reduced intensity and spatial extent in the same regions. In the ventral views (bottom), activation in the healthy group spans broadly across the posterior occipital cortices, while depressed participants again display markedly diminished engagement. This pattern points to altered visual processing network function in depression, consistent with prior reports of occipital cortical abnormalities, including disrupted dynamics at rest and altered connectivity with emotion-regulation systems Wu2023VisualCortexMDD , Qin2025OccipitalDynamicsDepression , Xie2024OccipitalPCCConnectivity .
Figure 5: 3D rendering of emotion predicted activation differences (Table 3) showing lateral (top) and ventral (bottom) views between healthy (left) and depressed (right) subjects. The color bar indicates normalized activation magnitudes, ranging from 0.07 (white) to 0.100 (red).
@@
Table 4: Multi-Trial Analysis Results: Original vs. Multi-Trial.
Brain Region Original (Imbalanced) Balanced Multi-Trial
H Mean (n=97) D Mean (n=37) p-val (MWU) Mean p Sig Trials (%) Effect Size (d)
@@
[…]
Hippocampus R 0.075 0.100 0.734 13.3 $2±\pm0$.15
@@
[…]
Orbitofrontal R 0.100 0.100 1.000 0.889 0.0 $1±\pm0$.08
[…]
@@
[…]
Note: Highlighted rows (shaded) indicate regions flagged as notable by our selection rule: either (a) $\geq 20\%$ of balanced multi-trial runs showed a significant difference (“Sig Trials”), or (b) mean effect size (Cohen’s $d$) >0.4>0.4. “Mean p” is the average p-value across 15 balanced trials. “H Mean” and “D Mean” are original-group means reported for reference. Missing MWU p-values are shown as “–” when not applicable.
The clustering quality ratio (depressed vs. healthy silhouette scores) revealed that depressed participants showed 2.2 - 2.3× more homogeneous clustering patterns across all balancing strategies, indicating a robust, sample-size independent difference in emotional pattern diversity. The analysis further revealed a previously unrecognized difference between healthy and depressed populations: depressed individuals demonstrate emotional pattern rigidity, a constraint in the diversity and flexibility of emotional responses koster2017cognitive . This finding aligns with emerging theories of depression emphasizing cognitive and behavioral inflexibility joormann2015cognitive , derubeis2017cognitive , and extends these concepts to emotional processing patterns derived from natural language. The clustering pattern differences are also consistent with neuroimaging findings showing reduced network flexibility in depression bassett2011dynamic , braun2020network , while the mapping of constrained emotional patterns to brain regions aligns with evidence of altered connectivity and reduced neural network switching in depressed individuals broyd2009default , kaiser2015large .
The high silhouette scores in the depressed group suggest stereotyped, predictable emotional expressions, while the variable clustering in healthy participants indicates adaptive emotional flexibility: the ability to express emotions across a broader range of patterns depending on context aldao2010emotion . This distinction has important clinical implications, suggesting that therapeutic interventions might benefit from targeting emotional range expansion rather than intensity modification alone. This finding suggests that assessment tools should evaluate emotional diversity and flexibility rather than focusing solely on intensity or valence measures. Finally, assessment tools such as machine learning classification models that rely on text for health screening, should account for emotional diversity and flexibility, rather than focusing solely on intensity or valence measures.
@@ 3.2 Experiment 2: Multiple Emotional States
The second experiment was performed on the GoEmotions dataset dataset_emotions , which includes 58,000 Reddit reddit comments manually labeled into 27 emotion categories (or neutral). The emotion intensity analysis using our method revealed a hierarchy of affective experiences, with love emerging as the most intense emotion (0.709), followed by joy (0.593) and relief (0.560). Negative emotions like sadness (0.486), fear (0.412), and anger (0.390) occupy middle-intensity positions. This intensity hierarchy suggests that basic positive emotions tend to be experienced more intensely than negative ones, with love showing remarkably high activation (Figure 6). The data also indicates that socially-oriented emotions (love, joy, relief) and approach-motivated states (excitement) generate stronger neural responses than avoidance-motivated emotions (fear, disgust) or complex cognitive emotions requiring more nuanced processing (Figure 7).
Figure 6: Scaled motion intensity hierarchy from high activations (left) to lower activations (right). Dark blue indicates positive emotions, with light blue indicating negative emotions.
@@
The emotion intensity hierarchy presented here reflects patterns derived from our lexicon-based scoring system and embedding-based clustering, mapped onto predefined brain coordinates. References to neuroimaging literature serve to contextualize rather than validate these patterns. The framework’s utility lies in generating testable predictions: for example, that texts expressing love would elicit stronger fMRI responses in regions our method associates with this emotion compared to texts expressing fear.
@@
These findings align with established emotion research, particularly regarding the valence-arousal relationship drevets2008 . Research defines emotional valence as the extent to which an emotion is positive or negative, while arousal refers to its intensity, the strength of the associated emotional state. The results supports the general principle that negative words tend to have higher arousal values and are perceived with higher intensity than positive words hastings2004 , while also showing positive emotions like love and joy to be at the top of the intensity scale.
@@
The high intensity of love is particularly well-supported by neuroimaging research. Meta-analyses have found that love recruits brain regions that mediate motivation, emotion, social cognition, and self-representation, including the ventral tegmental area, caudate nucleus, anterior cingulate gyrus, and middle frontal gyrus castanheira2019 . Further studies showed that positive emotions connect the prefrontal cortex to the nucleus accumbens, while negative emotions connect the nucleus accumbens to the amygdala brosch2022 , suggesting different neural pathways that could explain intensity differences.
@@
The positioning of joy as the second-highest intensity emotion is consistent with neuroscience research showing that the left prefrontal cortex is particularly associated with positive emotions including joy, with increased activity in the left prefrontal cortex correlated with positive emotional states davidson2004 . Research identifies positive emotions like happiness, interest, satisfaction, pride, and love as being generated by individuals in response to internal and external stimuli sliz2012 , supporting the results showing that these emotions cluster in the high-intensity range. The relatively low intensity of cognitive emotions aligns with research suggesting these require more complex processing okon2003 , but the moderate intensity of fear (0.412) is somewhat lower than might be expected given fear’s evolutionary importance phillips2003 .
To further assess the robustness and interpretability of the computational framework, three complementary validation analyses were conducted: i) quantification of PCA variance capture, ii) visualization of effect size distributions, and iii) comparison of emotion-intensity metrics to an established affective lexicon.
The cumulative variance explained by the 3D PCA reduction (Figure 2) was 8.98% of the original 1,536-dimensional embedding space, confirming that the reduced representation primarily supported spatial visualization rather than full variance preservation. To better capture nonlinear cluster relationships, a t-SNE projection of the full embedding was generated (Supplemental Figure S1), revealing clearer separations among emotion clusters consistent with those observed in the main analyses. To provide a more interpretable summary of group-level differences, the Cohen’s d values reported in Table 4 were visualized as a horizontal bar plot (Supplemental Figure S2). This plot highlighted consistent regional intensity differences between healthy and depressed groups, with the largest effects observed in the raphe nuclei and insula left regions.
To examine the validity of the custom emotion-intensity hierarchy (Figure 6), emotion clusters derived from Experiment 2 were compared against the Warriner et al. Valence–Arousal–Dominance (VAD) lexicon warriner2013norms . The cluster dominated by love (20.3%) and admiration (32.5%) exhibited the highest VAD Arousal (4.436) and custom intensity (0.483) scores, confirming convergence between the hierarchy and established affective measures.